The freezing point of water is 0.00°C at 1 atmosphere.  

A student dissolves 12.62 grams of aluminum acetate, Al(CH3COO)3 (204.1 g/mol), in 254.5 grams of water. Use the table of boiling and freezing point constants to answer the questions.

1. What is the molality of the solution?

2. What is the freezing point of the solution?

Solvent Formula Kb (°C/m) Kf (°C/m)
Water H2O 0.512 1.86
Ethanol CH3CH2OH 1.22 1.99
Chloroform CHCl3 3.67
Benzene C6H6 2.53 5.12
Diethyl ether CH3CH2OCH2CH3 2.02
Answer #1

1.

molality    = weight of solute*1000/gram molecular weight of solute* weight of solvent in g

                = 12.62*1000/(204.1*254.5)

                = 0.243m

Al(CH3COO)3 (aq) ----------> Al^3+ (aq) + 3CH3COO^- (aq)

i   = 4

Tf   = i*Kf*m

          = 4*1.86*0.243

          = 1.810C

Tf    = Tf solvent - Tf solution

1.81 = 0 -Tf solution

Tf solution = -1.810C

The freezing point of the solution    = -1.810C